OpenStudy (anonymous):

Graph 2x - 5y = 10 using the slope and y-intercept

OpenStudy (anonymous):

So in this one you need to make it into slope-intercept form, which is y=mx+b M=Slope B=y intercept

OpenStudy (anonymous):

So you want to get y on one side :) Does that make sense? :)

OpenStudy (anonymous):

So will that be 5y=-2y+10

OpenStudy (anonymous):

Is this right

OpenStudy (skullpatrol):

no

OpenStudy (anonymous):

-2/5x + 2

OpenStudy (anonymous):

How about this

OpenStudy (skullpatrol):

add 5y to both sides

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

what do you get?

OpenStudy (anonymous):

-5/2+ 2

OpenStudy (skullpatrol):

2x - 5y = 10 2x - 5y +5y = 10 +5y

OpenStudy (skullpatrol):

2x = 10 + 5y

OpenStudy (anonymous):

5 + 5/2

OpenStudy (skullpatrol):

2x - 10 = 10 + 5y -10 (subtract 10 from both sides)

OpenStudy (skullpatrol):

2x - 10 = 5y

OpenStudy (skullpatrol):

follow so far?

OpenStudy (anonymous):

yeah

OpenStudy (skullpatrol):

2x - 10 = 5y (1/5)(2x - 10) = (1/5)(5y) multiply both sides by 1/5

OpenStudy (skullpatrol):

(2/5)x - 2 = y

OpenStudy (skullpatrol):

y = mx + b

OpenStudy (skullpatrol):

got it?

OpenStudy (anonymous):

add 2

OpenStudy (skullpatrol):

why "add 2" ?

OpenStudy (skullpatrol):

y = mx + b y = (2/5)x - 2

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

m = 2/5, b = -2

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

Thanks for your help